OpenStudy (anonymous):

a similar situation is only an interior angle of a regular polygon measure 156 degrees. how many sides does this polygon have? i really dont understand this one...

OpenStudy (whpalmer4):

Okay, we use the same formula to find the sum of the angles, but then we set that equal to another expression that gives us the angle of the polygon given the number of sides. The angle is only dependent on the number of sides, and so is the the sum of the angles, so we have enough information to solve. \[{Sum}=(n-2)180 \]\[{Sum}/156 = n\]\[{Sum}=156n\]Substitute one equation into the other\[156n=(n-2)180\]\[156n=180n-360\]Solve that for n and you've got your answer. Bonus points if you've ever heard of a pentadecagon :-)

OpenStudy (amoodarya):

 do you know sum of angle? if ok divide it to n